OpenStudy (anonymous):

Sue bought five packages of cookies and two sacks of potato chips for $4.75. Later she bought three more packages of cookies and two additional sacks of potato chips for $2.89. Find the price of a package of cookies.

OpenStudy (anonymous):

c = cookie p = potato chips 5c + 2p = 4.75 3c + 2p = 2.89 subtract second equation from first 2c = 1.96 c = 0.98

OpenStudy (radar):

Check the subtraction 4.75-2.89=??????